brother 3 yrs ago was 2/3 sisters age.in one year he will be 3/4 her age. find his and her age.
Answer by stanbon(75887)

brother 3 yrs ago was 2/3 sisters age.in one year he will be 3/4 her age. find his and her age.
===========
b-3 = (2/3)(s-3)
b+1 = (3/4))(s+1)
---------------------
3b-9 = 2s-6
4b+4 = 3s+3
----------------------
3b - 2s = 3
4b - 3s = -1
-----
12b - 8s = 12
12b - 9s = -3
-----
Subtract and solve for "s"::
s = 15 (sister's age now)
Solve for "b"::
3b - 2s = 3
3b - 30 = 3
b = 11 years (brother's age now)
-------------
Cheers,
Stan H.
-----------